# Less is more? Rewards in RL for Cyber Defence

### v1

@@ Abstract
The last few years has seen an explosion of interest in autonomous cyber defence agents based on deep reinforcement learning. Such agents are typically trained in a cyber gym environment, also known as a cyber simulator, at least 32 of which have already been built. Most, if not all cyber gyms provide dense “scaffolded” reward functions which combine many penalties or incentives for a range of (un)desirable states and costly actions. Whilst dense rewards help alleviate the challenge of exploring complex environments, yielding seemingly effective strategies from relatively few environment steps; they are also known to bias the solutions an agent can find, potentially towards suboptimal solutions. This is especially a problem in complex cyber environments where policy weaknesses may not be noticed until exploited by an adversary. Sparse rewards could offer preferable or more effective solutions and have been overlooked by cyber gyms to date. In this work we set out to evaluate whether sparse reward functions might enable training more effective cyber defence agents. Towards this goal we first break down several evaluation limitations in existing work by proposing a ground truth evaluation score that goes beyond the standard RL paradigm used to train and evaluate agents. By adapting a well-established cyber gym to accommodate our methodology and ground truth score, we propose and evaluate two sparse reward mechanisms and compare them with a typical dense reward. Our evaluation considers a range of network sizes, from 2 to 50 nodes, and both reactive and proactive defensive actions. Our results show that sparse rewards, particularly positive reinforcement for an uncompromised network state, enable the training of more effective cyber defence agents. Furthermore, we show that sparse rewards provide more stable training than dense rewards, and that both effectiveness and training stability are robust to a variety of cyber environment considerations.
@@ 2.1 Yawning Titan
Yawning Titan (YT) is a well-established cyber gym providing an abstract, graph based network simulation environment for training defensive ACD agents  [1]. YT is highly configurable, allowing for customisable game dynamics, actions and reward functions. We configure YT to simulate a linear network structure with a single fixed entry node for the red agent. The attacker can only spread from an already compromised node, moving laterally through the network. The goal of the blue agent is to defend the network by minimising the number of compromised nodes. Conversely, the red agent aims to compromise as many nodes as possible.
@@ 3 Results and Discussion
Firstly, we examine the impact of each reward function using the basic action space and the standard agent order of red then blue. We configure the red agent with a 90% probability of performing a basic attack (which always succeeds) each time step, alternatively laying dormant i.e., does nothing 10% of the time. Consequently, the best possible evaluation score for any defensive agent is 0.90.9-0.9- 0.9. This would mean the red agent is making no progress through the network and is confined to, at most, a single node. Concerning $DV\bar$$\text${DV}}over¯ start_ARG DV end_ARG across time (i.e., training reliability) a score of 0 is the best-case and indicates minimal fluctuations in the de-trended performance throughout training. As shown in Table 2, only the sparse positive reward achieves an optimal average evaluation score for all network sizes. The negative and dense reward functions show significantly worse evaluation scores and also decrease in performance as the network size increases. Only in the smallest network size of 2 nodes does the dense reward achieve an optimal score. Considering variability during training, both the positive and negative sparse rewards show no significant de-trended performance fluctuations throughout training. Taken together, the positive sparse reward alone shows the combination of both high training reliability and optimal final performance.
@@
Next, we consider the impact of agent order by maintaining the basic action space but with the alternative agent order of blue then red. As the red agent attack rate is unchanged from 90%, the optimal evaluation score remains 0.90.9-0.9- 0.9 on average and the best-case $DV\bar$$\text${DV}}over¯ start_ARG DV end_ARG is still $0$. The results in Table 3 show that this setting increases the difficulty of network defence irrespective of the reward function. Only the sparse positive reward achieves the optimal average evaluation score for the smallest network size of 2 nodes. For networks with 5 nodes, the sparse positive reward also provides the best performance. Thereafter, performance decreases and the dense reward shows superior, although sub-optimal, average scores. Compared with the red then blue agent order, $DV\bar$$\text${DV}}over¯ start_ARG DV end_ARG across time scores are greatly elevated. This is most severely the case for the dense reward which increases with the number of network nodes. The sparse positive reward is uniquely antagonised by the blue then red agent order as the only circumstances for incentive, a network with no compromised nodes, is dependent on the red agent not performing a basic attack. As this occurs with only a 10% probability, its rewards are made particularly sparse by this configuration.
@@
Lastly, Figure 1 compares the episodic rewards and evaluation score throughout training when the network size is 5, the agent order is red then blue, and agents use the basic action space. Despite offering comparatively infrequent incentives, the sparse positive reward shows both a steeper policy-improvement gradient and greater stability across time.

### v2

@@ Abstract
The last few years have seen an explosion of interest in autonomous cyber defence agents based on deep reinforcement learning. Such agents are typically trained in a cyber gym environment, also known as a cyber simulator, at least 32 of which have already been built. Most, if not all cyber gyms provide dense “scaffolded” reward functions which combine many penalties or incentives for a range of (un)desirable states and costly actions. Whilst dense rewards help alleviate the challenge of exploring complex environments, yielding seemingly effective strategies from relatively few environment steps; they are also known to bias the solutions an agent can find, potentially towards suboptimal solutions. This is especially a problem in complex cyber environments where policy weaknesses may not be noticed until exploited by an adversary. Sparse rewards could offer preferable or more effective solutions and have been overlooked by cyber gyms to date. In this work we set out to evaluate whether sparse reward functions might enable training more effective cyber defence agents. Towards this goal we first break down several evaluation limitations in existing work by proposing a ground truth evaluation score that goes beyond the standard RL paradigm used to train and evaluate agents. By adapting a well-established cyber gym to accommodate our methodology and ground truth score, we propose and evaluate two sparse reward mechanisms and compare them with a typical dense reward. Our evaluation considers a range of network sizes, from 2 to 50 nodes, and both reactive and proactive defensive actions. Our results show that sparse rewards, particularly positive reinforcement for an uncompromised network state, enable the training of more effective cyber defence agents. Furthermore, we show that sparse rewards provide more stable training than dense rewards, and that both effectiveness and training stability are robust to a variety of cyber environment considerations.
@@ 2.1 Yawning Titan
Yawning Titan (YT) is a well-established cyber gym providing an abstract, graph-based network simulation environment for training defensive ACD agents  [1]. YT is highly configurable, allowing for customisable game dynamics, actions and reward functions. We configure YT to simulate a linear network structure with a single fixed entry node for the red agent. The attacker can only spread from an already compromised node, moving laterally through the network. The goal of the blue agent is to defend the network by minimising the number of compromised nodes. Conversely, the red agent aims to compromise as many nodes as possible.
@@ 3 Results and Discussion
Firstly, we examine the impact of each reward function using the basic action space and the standard agent order of red then blue. We configure the red agent with a 90% probability of performing a basic attack (which always succeeds) each time step, alternatively laying dormant i.e., does nothing 10% of the time. Consequently, the best possible evaluation score for any defensive agent is 0.90.9-0.9- 0.9. This would mean the red agent is making no progress through the network and is confined to, at most, a single node. Concerning $DV\bar$$\text${DV}}over¯ start_ARG DV end_ARG across time (i.e., training reliability) a score of 0 is the best-case and indicates minimal fluctuations in the de-trended performance throughout training. As shown in Table 2, only the sparse positive reward achieves an optimal average evaluation score for all network sizes. The negative and dense reward functions show significantly worse evaluation scores and also decrease in performance as the network size increases. The dense reward achieves an optimal score only in the smallest network size of 2 nodes. Considering variability during training, neither the positive nor the negative sparse rewards show significant de-trended performance fluctuations throughout training. Taken together, the positive sparse reward alone shows the combination of both high training reliability and optimal final performance.
@@
Next, we consider the impact of agent order by maintaining the basic action space but with the alternative agent order of blue then red. As the red agent attack rate is unchanged from 90%, the optimal evaluation score remains 0.90.9-0.9- 0.9 on average and the best-case $DV\bar$$\text${DV}}over¯ start_ARG DV end_ARG is still $0$. The results in Table 3 show that this setting increases the difficulty of network defence irrespective of the reward function. Only the sparse positive reward achieves the optimal average evaluation score for the smallest network size of 2 nodes. For networks with 5 nodes, the sparse positive reward also provides the best performance. Thereafter, performance decreases and the dense reward shows superior, although sub-optimal, average scores. Compared with the red then blue agent order, $DV\bar$$\text${DV}}over¯ start_ARG DV end_ARG across time scores are greatly elevated. This is most severely the case for the dense reward which increases with the number of network nodes. The sparse positive reward is uniquely antagonised by the blue then red agent order as the only circumstances for incentive, a network with no compromised nodes, is dependent on the red agent not performing a basic attack. As this occurs with only a 10% probability, sparse positive rewards are made particularly sparse by this configuration.
@@
Figure 1 compares the episodic rewards and evaluation score throughout training when the network size is 5, the agent order is red then blue, and agents use the basic action space. Despite offering comparatively infrequent incentives, the sparse positive reward shows both a steeper policy-improvement gradient and greater stability across time.